# Why Steering Works: Toward a Unified View of Language Model Parameter Dynamics

### v1

@@ Abstract
Methods for controlling large language models (LLMs), including local weight fine-tuning, LoRA-based adaptation, and activation-based interventions, are often studied in isolation, obscuring their connections and making comparison difficult. In this work, we present a unified view that frames these interventions as dynamic weight updates induced by a control signal, placing them within a single conceptual framework. Building on this view, we propose a unified preference-utility analysis that separates control effects into preference, defined as the tendency toward a target concept, and utility, defined as coherent and task-valid generation, and measures both on a shared log-odds scale using polarity-paired contrastive examples. Across methods, we observe a consistent trade-off between preference and utility: stronger control increases preference while predictably reducing utility. We further explain this behavior through an activation manifold perspective, in which control shifts representations along target-concept directions to enhance preference, while utility declines primarily when interventions push representations off the model’s valid-generation manifold. Finally, we introduce a new steering approach SPLIT guided by this analysis that improves preference while better preserving utility.
@@ 2.1 Intermediate Representations in LLMs
For example, in an FFN block, the up-projection is computed as $\mathbf{h}_{\mathrm{mid}}=\mathbf{W}_{\mathrm{up}}\mathbf{h}_{\mathrm{in}}+\mathbf{b}_{\mathrm{up}}$, followed by a non-linear activation, $\mathbf{h}_{\mathrm{mid,act}}=\sigma(\mathbf{h}_{\mathrm{mid}})$, and then the down-projection is computed as $\mathbf{h}_{\mathrm{out}}=\mathbf{W}_{\mathrm{down}}\,\mathbf{h}_{\mathrm{mid,act}}+\mathbf{b}_{\mathrm{down}}$.
Similarly, the $Q$, $K$, $V$, and output projections in the attention module follow the same affine form as in Eq. 2.
@@ 3 Unified View of Dynamic Weights in Inference
We present a unified framework for dynamic interventions during inference. Our unified view has three components: (i) a unified measurement view based on preference/utility log-odds, (ii) a unified dynamic weights intervention view that expresses local weight updates, LoRA, and activation steering as dynamic weight updates, and (iii) a unified dynamics observation showing consistent preference–utility response patterns across intervention forms.
@@ 3.1 Unified Analysis View: Preference and Utility Log-Odds
3.2 Unified Dynamic Weight Formulation
We propose a unified framework that encompasses both parameter update methods and activation steering methods, by viewing them as dynamic weight update. In this formulation, both classes of methods can be expressed under a shared affine transformation view of intermediate representations (see Section 2).
@@ 5.1 Preference–Utility Joint Optimization
Building on the preceding mechanistic analysis, we propose Steering with Preference–UtiLity IntervenTion (SPLIT), a training objective that improves preference while delaying utility degradation by extending the effective linear regime of activation interventions.
@@ 6 Related Work
Most activation steering methods assume linear structure in activation space, controlling concepts by adding scaled direction vectors to hidden states (Mikolov et al., 2013; Pennington et al., 2014; Nanda et al., 2023; Tigges et al., 2023; Park et al., 2024; Wang et al., 2024; Yao et al., 2025; zhang2026locate,hu2026towards). Building on this view, Bigelow et al. (2025) show that steering yields an approximately linear trend in posterior odds, but mainly in the small-scale regime. Recent studies further report non-monotonic or adverse effects under stronger steering, challenging a naive global linearity assumption (Bricken et al., 2023; Wollschläger et al., 2025). Meanwhile, representation-manifold work provides a complementary geometric lens for understanding steering and its limitations (Modell et al., 2025; Li and He, 2025; Xie et al., 2025).

### v2

@@ Abstract
Methods for controlling large language models (LLMs), including local weight fine-tuning, LoRA-based adaptation, and activation-based interventions, are often studied in isolation, obscuring their connections and making comparison difficult. In this work, we present a unified view that frames these interventions as dynamic weight updates induced by a control signal, placing them within a single conceptual framework. Building on this view, we propose a unified preference-utility analysis. This analysis separates control effects into two components: preference, defined as the tendency toward a target concept, and utility, defined as coherent and task-valid generation. Both components are measured on a shared log-odds scale using polarity-paired contrastive examples. Across methods, we observe a consistent trade-off between preference and utility: stronger control increases preference while predictably reducing utility. We further explain this behavior through an activation manifold perspective, in which control shifts representations along target-concept directions to enhance preference, while utility declines primarily when interventions push representations off the model’s valid-generation manifold. Finally, we introduce a new steering approach SPLIT guided by this analysis that improves preference while better preserving utility [^1] https://github.com/zjunlp/EasyEdit/blob/main/examples/SPLIT.md..
@@ 2.1 Intermediate Representations in LLMs
For example, in an FFN block, the up-projection is computed as $\mathbf{h}_{\mathrm{mid}}=\mathbf{W}_{\mathrm{up}}\mathbf{h}_{\mathrm{in}}+\mathbf{b}_{\mathrm{up}}$, followed by a non-linear activation, $\mathbf{h}_{\mathrm{mid,act}}=\sigma(\mathbf{h}_{\mathrm{mid}})$, and then the down-projection is computed as $\mathbf{h}_{\mathrm{out}}=\mathbf{W}_{\mathrm{down}}\,\mathbf{h}_{\mathrm{mid,act}}+\mathbf{b}_{\mathrm{down}}$. Similarly, the $Q$, $K$, $V$, and output projections in the attention module follow the same affine form as in Eq. 2.
@@ 3 Unified View of Dynamic Weights in Inference
We present a unified framework for dynamic interventions during inference, as illustrated in Figure 1. Our unified view has three components: (i) a unified dynamic weights intervention view that expresses local weight updates, LoRA, and activation steering as dynamic weight updates, (ii) a unified measurement view based on preference/utility log-odds, and (iii) a unified dynamics observation showing consistent preference–utility response patterns across intervention forms.
@@
3.1 Unified Dynamic Weight Formulation
We propose a unified framework that encompasses both parameter update methods and activation steering methods, by viewing them as dynamic weight updates. Under this formulation, both can be expressed through a shared affine transformation view of intermediate representations; detailed derivations and formulations are provided in Section 2.
@@ 5.1 Preference–Utility Joint Optimization
Building on the preceding mechanistic analysis, we propose Steering with Preference–UtiLity IntervenTion (SPLIT), a training objective improving preference while delaying utility degradation.
@@ 6 Related Work
Most activation steering methods assume linear structure in activation space, controlling concepts by adding scaled direction vectors to hidden states (Mikolov et al., 2013; Pennington et al., 2014; Nanda et al., 2023; Tigges et al., 2023; Park et al., 2024; Wang et al., 2024; Yao et al., 2025; Zhang et al., 2026; Hu et al., 2026). Building on this view, Bigelow et al. (2025) show that steering yields an approximately linear trend in posterior odds, but mainly in the small-scale regime. Recent studies further report non-monotonic or adverse effects under stronger steering, challenging a naive global linearity assumption (Bricken et al., 2023; Wollschläger et al., 2025). Meanwhile, representation-manifold work provides a complementary geometric lens for understanding steering and its limitations (Modell et al., 2025; Li and He, 2025; Xie et al., 2025).